OpenStudy (anonymous):

Which property is used to simplify the following expression? 2003-01-05-00-00_files/i0270000.jpg A. commutative property B. distributive property C. associative property D. inverse property

OpenStudy (unklerhaukus):

the .jpeg isn't a link

OpenStudy (unklerhaukus):

Commutative Property\[a\star b=b\star a\] Distribution Property\[a\star(b+ c)=a\star b+ a\star c\] Associative property\[a\star(b\star c)=(a\star b)\star c\] Additive inverse property \[a+(-a)=0\] Multiplicative inverse property \[a\times(\frac1a)=1\]